OpenStudy (vishweshshrimali5):

The cartesian equations of a line are x = ay + b, z = cy + d. Find its direction ratios and reduce it to vector form.

OpenStudy (jamesj):

By way of example, consider the line y = 2x + 3 then for every +1 of x, y changes by +2 Hence in vector form, the line is (1,2)t + (0,3)

OpenStudy (jamesj):

Apply this idea to your problems.

OpenStudy (vishweshshrimali5):

sorry sir that is not working. Will you please explain me how to solve it

OpenStudy (jamesj):

Ok, so what is the vector form of the line y = x ?

OpenStudy (vishweshshrimali5):

\[y = x => y - x = 0 => j - i = 0 => i = j\] I hope I am right

OpenStudy (jamesj):

Well, not quite, because i is not j. But what is true is that for any t, the vector t i + t j is on the line.

OpenStudy (jamesj):

This is expression is what I mean by the vector form of the line. We can also write it as v = t(1,1) or v = t (i + j)